OpenStudy (anonymous):

Solve x2 + 8x − 3 = 0 using the completing-the-square method. A.x = four plus or minus the square root of three B.x = negative four plus or minus the square root of three C.x = four plus or minus the square root of nineteen D.x = negative four plus or minus the square root of nineteen

OpenStudy (welshfella):

(x + 4)^2 - 16 - 3 = 0 can you see why i subtract 16 ?

OpenStudy (anonymous):

nope lol

OpenStudy (welshfella):

ok so to complete the square you first divide the coefficient of x in the original equation by 2 so here its +8 / 2 = +4 so we write (x + 4)^2 - thats the square bit but if you expand this it comes to x^2 + 8x + 16. So to make it equal to x^2 + 8x we have to subtract 16

OpenStudy (anonymous):

okay i dont undertand how there is a squaare root in the answer tho

OpenStudy (welshfella):

as an identity in general form its is x^2 + bx = (x + (b/2)^2 - ( b/2)^2

OpenStudy (welshfella):

because we have a square in the equation (x + 4)^2 -3 - 16 = 0 (x + 4)^2 = 19 taking square roots x + 4 = +/- sqrt 19

OpenStudy (welshfella):

x = -4 +/- sqrt 19